# Analysis of Josephson Junction Barrier Variation - a Combined STEM, Breakdown and Monte-Carlo Approach

### v1

@@ III Scanning Transmission Electron Microscopy
Figure 4: Experimental STEM-EDS imaging of an Al/AlOxAl barrier of the oxygen peak. (a-c) show heatmaps indicating the oxygen peak with the edges of the barrier as detected by Kernel integration overlaid to the maps. Different values of $\delta$ are used to define the kernel for the three maps (0, 0.2, 0.4 respectively). (d-f) show histograms of thicknesses as inferred from the different kernel integrations.
@@
STEM includes a family of techniques which allow for sub-nanometre resolution imaging of materials with elemental contrast which are natural candidates to investigate questions about the thickness distribution of the barrier. These techniques can be applied to JJs where the barrier has been capped with a top electrode, i.e. after the full fabrication process has taken place. We first consider a protocol to measure variations in thickness in the projected 2D image. We do this considering images of the oxygen content in a barrier collected by STEM energy-dispersive X-ray spectroscopy (EDS) mapping. Imaging is performed on a lamella which is fabricated by focused ion beam cutting and lift out of a $\sim$30 nm thick cross section of the JJ. In Fig. 4 we present an example image of the oxygen composition. The edge of the barrier is identified by kernel integration using OpenCV [41]. We define two kernels to identify the two edges where each kernel has an asymmetry factor $\delta$. We describe the construction of the integration kernels in the supplemental information. Given that the oxygen content in the barrier increases over a finite length, the edge of the barrier could be defined at several points. By changing $\delta$ we systematically alter where the kernel integration protocol identifies the edge. In Fig. 4 (a-c) we show the same image with edges identified using 0, 0.2, 0.4 as values of $\delta$ respectively. In analysis we explored larger values of $\delta$ and found the edges were obviously within the junction, so care must be taken in selecting this asymmetry value.
@@
We measure the thickness of the barrier by finding the distance between the two edges for each column of pixels. The average of the thickness from these different kernels matches previous measurements by TEM well giving an average thickness of $\sim$ 1.9 nm - 2.4 nm depending on the kernel [23]. We can also compute the distribution of thicknesses shown in the histograms Fig. 4 (d-f) and fit each of them both to normal and log-normal distributions. We routinely see that the log-normal distribution fits the thickness distribution better than the normal distribution. In the literature the results of this type of analysis have been interpreted as the statistical distribution of thicknesses inside a barrier, a claim we are explicitly not making here.
@@
Arbitrary choices, such as the asymmetry factor in the kernels, can change what we infer about the barrier. We attribute this both to (i) details of the barrier including such as the oxygen content likely changing over a finite distance and the non-uniform barrier thickness and (ii) the STEM image being a projection which averages the barrier. As the barrier is grown on a bottom electrode which isn’t perfectly flat the imaging beam is at times intersecting both electrode and barrier. We wish to understand (i) from techniques such as STEM imaging as they will determine the behaviour of a JJ. However (ii) represents difficulties posed in trying to identify thickness variations in this material system by STEM. We suggest that we should report 1.9 nm $<\langle t\rangle<$ 2.4 nm, i.e. the variation in thickness implied by different kernels is what defines the error bars on the measurement of average barrier thickness.
@@
Figure 5: (a, b) two high resolution AFM scans of representative bottom leads (region indicated in Fig. 1 a) In (a, b) five 30 $\times$ 100 nm areas are marked by dashed lines. These regions are used to simulate STEM profiles shown in (c-g) assuming a 2 nm oxide barrier coating the AFM topography with the labels matching the AFM region to the STEM profile inserted into the boxes. (c-g) Edges of simulated oxygen content are detected using a symmetric kernels and shown overlaid, as well as measurements of average thickness and thickness standard deviations being noted in each figure. The labels (c - g) are in the same side of the scan region in the simulated STEM profile and the AFM regions.
@@
We generate a series of barriers assuming that the thickness of the barrier is uniformly 2 nm across the junction. Using a uniform barrier thickness means that any thickness variation inferred from the simulated STEM image arises from (iii) and (iv) above, i.e. experimental nonidealities which could change what we infer from analysis of STEM images. We draw the bottom electrode topography from AFM measurements shown in Fig. 5 (a,b).  [^3] Noise at the $\sim$0.6 nm range is present in some regions of these AFM scans, which we found hard to exclude from measurements. This represents $\sim$10 % of the total topographic variation in these AFM maps and is not present across the whole scans and therefore we believe not a dominant contributor to the following analysis.. We infer the thickness of the barriers from the simulated STEM images (Fig. 5 c-g) using symmetric kernels ($\delta=0$) as described for Fig. 4 adjusting kernel sizes to account for the different pixel sizes. As expected, a perfectly uniform barrier appears to have varying thickness when viewed in projection due to roughness in the underlying barrier. We also find that the thickness values and thickness standard deviation change depending which AFM area is used to build the projection. All but one average thickness measurements are within 10 % of the barrier thickness input to the simulation, although the standard deviation of barrier thickness varies by a factor of 8$\sim$ 8$\times$ 8 ×. This suggests that STEM is a good technique to measure the average thickness of a barrier (reporting errors as described above) but calls into question the robustness of using STEM profiles to infer the thickness distribution in a junction. It also shows that multiple areas should be measured and inferring average thickness values from single STEM images may introduce errors. We therefore include additional STEM images with analogous processing to that performed in Fig. 4 in the supplemental material. These suggest similar thicknesses, but also include regions of topographic features which slightly increase the inferred average thickness.
@@
Sometimes it is straightforward to identify regions where apparent thickness broadening occurs due to topographic variation in the bottom electrode (e.g. RHS of Fig. 5 c) where it’s possible to resolve two barriers superimposed on one another. Here the kernel edge detection gets confused shown by the discontinuities in the edge it identifies. In the supplemental material we show an example of experimental STEM imaging with these types of features, compared to a similar STEM simulation. In Fig. 5 (d) there is also a topographic gradient along the projection axis, but, in this instance, the topographic change is smaller in magnitude and so in the simulated STEM image it appears only as a broadening of the barrier. It is therefore not straightforwardly possible to systematically distinguish apparent broadening from topography variation and true barrier variation, although it is possible to exclude areas where the topographic broadening effects are largest.
@@
Comparing experimental (Fig. 4) and simulated (Fig. 5) STEM images, barriers appear rougher in the experimental STEM data than the simulated STEM. The topography of the electrode used to simulate STEM images may be artificially smoothed due to the AFM tip radius of 2 nm. This would mean that features of the size of apparent roughness seen in Fig. 4 would not be apparent with this AFM tip. The barriers we measure here are also oxidised which can change the topography of the surface. While there may be features sufficiently small that they are not resolved by AFM, both AFM maps in Fig. 5 have topographic features are large relative to the AFM resolution, and as we show, couple importantly into the inferred STEM projections. Using STM may allow higher resolution topographic maps to perform more robust inferences of STEM profiles, and can also be performed in situ before barrier oxidation modifies the topography. These more detailed maps might find uses, for instance in reverse Monte-Carlo approaches [43] to extract quantitative information about the barrier thickness and thickness distributions from STEM.
@@ IV Conclusions
0.73 $\pm$ 0.02 2.10 $\pm$ 0.55 1.45$\pm$ 0.38 1.02 $\pm$ 0.21 2.39 $\pm$ 0.32 2.13 $\pm$ 0.22 1.94 $\pm$ 0.17
Table 1: A summary of different barrier thicknesses extracted throughout this manuscript. Fitting the Simmons model returns a single-valued thickness and the error represents the standard deviation from fitting 598 IVs. Monte-Carlo simulations give thickness by fixing the area of a junction and allowing the barrier thickness to vary. The thickness value is the average thickness whereas the errors represent a genuine thickness distribution in the barrier. STEM-EDS measurements are direct measurements of thickness by atomic scale microscopy. As shown here, the thickness of the barrier is sensitive to image processing (i.e. choice of $\delta$). Errors here represent standard deviations of inferred average thickness for a specific choice of $\delta$ but do not represent thickness distributions. Without an un-biased approach to image processing, the true error on thickness measurements is likely the variation in thickness between different values of $\delta$.
In this article we have aimed to understand oxide barriers in JJs due to their technological importance in superconducting qubits. A summary of different thicknesses inferred within this paper is shown in Table 1. We fit the Simmons model to non-linear IVs to extract a single-valued thickness value to each junction $\sim$0.75 nm. We build Monte-Carlo protocols to simulate the IV properties of barriers with multi-valued thicknesses drawn from example statistical distributions Here we find that the distribution of thicknesses present in the junction, results in IVs that, when fitted, would imply a single-value thickness smaller than the average thickness. This is because conductivity is exponentially dependent on the barrier thickness, so any thinner-than-average regions contribute more than similarly thicker-than-average regions. These simulations indicate a barrier that is $\sim$1 nm thick with a standard deviation of $\sim$0.2 nm thick, however using different barrier heights, we infer different thickness distributions. The rectangular potential barrier used in the Simmons model likely results in an under-estimate of thickness relative to more realistic, smoothly increasing barriers. By considering the Monte Carlo simulations relative to STEM measurements we constrain the barrier height to the range 0.8 eV ϕless-than-or-similar-toabsentitalic-$\lesssim\phi\lesssim$ 1.22 eV.

### v2

@@ III Scanning Transmission Electron Microscopy
Figure 4: Experimental STEM-EDS imaging of an Al/AlOx/Al barrier of the oxygen peak. (a-c) show heatmaps indicating the oxygen peak with the edges of the barrier as detected by Kernel integration overlaid to the maps. Different values of $\delta$ are used to define the kernel for the three maps (0, 0.2, 0.4 respectively). (d-f) show histograms of thicknesses as inferred from the different kernel integrations.
@@
STEM includes a family of techniques which allow for sub-nanometre resolution imaging of materials with elemental contrast which are natural candidates to investigate questions about the thickness distribution of the barrier. These techniques can be applied to JJs where the barrier has been capped with a top electrode, i.e. after the full fabrication process has taken place. We first consider a protocol to measure variations in thickness in the projected 2D image. We do this considering images of the oxygen content in a barrier collected by STEM energy-dispersive X-ray spectroscopy (EDS) mapping. Imaging is performed on a lamella which is fabricated by focused ion beam cutting and lift out of a $\sim$30 nm thick cross section of the JJ. In Fig. 4 we present an example image of the oxygen composition. The edge of the barrier is identified by kernel integration using OpenCV [45]. We define two kernels to identify the two edges where each kernel has an asymmetry factor $\delta$, which differentially changes the weight of pixels on the left and right hand side of the kernel. We describe the construction of the integration kernels in the supplemental information. Given that the oxygen content in the barrier increases over a finite length, the edge of the barrier could be defined at several points. By changing $\delta$ we systematically alter where the kernel integration protocol identifies the edge. In Fig. 4 (a-c) we show the same image with edges identified using 0, 0.2, 0.4 as values of $\delta$ respectively. In analysis we explored larger values of $\delta$ and found the edges were obviously within the junction, so care must be taken in selecting this asymmetry value.
@@
We measure the thickness of the barrier by finding the distance between the two edges for each column of pixels. The average of the thickness from these different kernels matches previous measurements by TEM well giving an average thickness of $\sim$ 1.9 nm - 2.4 nm depending on the kernel [26]. We can also compute the distribution of thicknesses shown in the histograms Fig. 4 (d-f) and fit each of them both to normal and log-normal distributions. We routinely see that the log-normal distribution fits the thickness distribution better than the normal distribution. In the literature the results of this type of analysis have been interpreted as the statistical distribution of thicknesses of a barrier, a claim we are explicitly not making here.
@@
Arbitrary choices, such as the asymmetry factor in the kernels, can change what we infer about the barrier. We attribute this both to (i) details of the barrier including such as the oxygen content likely changing over a finite distance and the non-uniform barrier thickness and (ii) the STEM image being a projection which averages the barrier. As the barrier is grown on a bottom electrode which isn’t perfectly flat the imaging beam is at times intersecting both electrode and barrier. We wish to understand (i) from techniques such as STEM imaging as they will determine the behaviour of a JJ. However (ii) represents difficulties posed in trying to identify thickness variations in this material system by STEM. We suggest that we should report 1.9 nm $\langle t\rangle<$ 2.4 nm, i.e. the variation in thickness implied by different kernels is what defines the error bars on the measurement of average barrier thickness.
@@
Figure 5: (a, b) two high resolution AFM scans of representative bottom leads (region indicated in Fig. 1 a). The route mean squared roughness for the areas shown are 0.7 and 0.5 nm respectively. In (a, b) five 30 $\times$ 100 nm areas are marked by dashed lines. These regions are used to simulate STEM profiles shown in (c-g) assuming a 2 nm oxide barrier coating the AFM topography with the labels matching the AFM region to the STEM profile inserted into the boxes. (c-g) Edges of simulated oxygen content are detected using a symmetric kernels and shown overlaid, as well as measurements of average thickness and thickness standard deviations being noted in each figure. The labels (c - g) are in the same side of the scan region in the simulated STEM profile and the AFM regions.
@@
We generate a series of barriers assuming that the thickness of the barrier is uniformly 2 nm across the junction. Using a uniform barrier thickness means that any thickness variation inferred from the simulated STEM image is fictitious, indicative of what we may expect to see from experimental nonidealities which could change what we infer from analysis of STEM images. We draw the bottom electrode topography from AFM measurements of the bottom lead shown in Fig. 5 (a,b). These scans were taken from a region approximately overlapping the text ‘Bottom Lead’ in Fig. 1 (a).  [^2] Noise at the $\sim$0.6 nm range is present in some regions of these AFM scans, which we found hard to exclude from measurements. This represents $\sim$10 % of the total topographic variation in these AFM maps and is not present across the whole scans and therefore we believe not a dominant contributor to the following analysis.. We infer the thickness of the barriers from the simulated STEM images (Fig. 5 c-g) using symmetric kernels ($\delta=0$) as described for Fig. 4 adjusting kernel sizes to account for the different pixel sizes. As expected, a perfectly uniform barrier appears to have varying thickness when viewed in projection due to roughness in the underlying barrier. We also find that the thickness values and thickness standard deviation change depending which AFM area is used to build the projection. All but one average thickness measurements are within 10 % of the barrier thickness input to the simulation, although the standard deviation of barrier thickness varies by a factor of $\sim 7\times$. This suggests that STEM is a good technique to measure the average thickness of a barrier (reporting errors as described above) but calls into question the robustness of using STEM profiles to infer the thickness distribution in a junction. It also shows that multiple areas should be measured and inferring average thickness values from single STEM images may introduce errors. We therefore include additional STEM images with analogous processing to that performed in Fig. 4 in the supplemental material. These suggest similar thicknesses, but also include regions of topographic features which slightly increase the inferred average thickness.
@@
Sometimes it is straightforward to identify regions where apparent thickness broadening occurs due to topographic variation in the bottom electrode (e.g. RHS of Fig. 5 c) where it’s possible to resolve two barriers superimposed on one another. Here the kernel edge detection does not work well, shown by the discontinuities in the edge it identifies. In the supplemental material we show an example of experimental STEM imaging with these types of features, compared to a similar STEM simulation. In Fig. 5 (d) there is also a topographic gradient along the projection axis, but, in this instance, the topographic change is smaller in magnitude and so in the simulated STEM image it appears only as a broadening of the barrier. It is therefore not straightforwardly possible to systematically distinguish apparent broadening from topography variation and true barrier variation, although it is possible to exclude areas where the topographic broadening effects are largest.
@@
Comparing experimental (Fig. 4) and simulated (Fig. 5) STEM images, barriers appear rougher in the experimental STEM data than the simulated STEM. The topography of the electrode used to simulate STEM images may be artificially smoothed due to the AFM tip radius of 2 nm. This would mean that features of the size of apparent roughness seen in Fig. 4 would not be apparent with this AFM tip. The barriers we measure here are also oxidised which can change the topography of the surface. While there may be features sufficiently small that they are not resolved by AFM, both AFM maps in Fig. 5 have topographic features that are large relative to the AFM resolution, and as we show, couple importantly into the inferred STEM projections. Using STM may allow higher resolution topographic maps to perform more robust inferences of STEM profiles, and can also be performed in situ before barrier oxidation modifies the topography. These more detailed maps might find uses, for instance in reverse Monte-Carlo approaches [47] to extract quantitative information about the barrier thickness and thickness distributions from STEM.
@@ IV Conclusions
[…]
Table 1: A summary of different barrier thicknesses extracted throughout this manuscript. Fitting the Simmons model returns a single-valued thickness and the error represents the standard deviation from fitting 597 IVs. Monte-Carlo simulations give thickness by fixing the area of a junction and allowing the barrier thickness to vary. The thickness value is the average thickness whereas the errors represent a genuine thickness distribution in the barrier. STEM-EDS measurements are direct measurements of thickness by atomic scale microscopy. As shown here, the thickness of the barrier is sensitive to image processing (i.e. choice of $\delta$). Errors here represent standard deviations of inferred average thickness for a specific choice of $\delta$ but do not represent thickness distributions. Without an un-biased approach to image processing, the true error on thickness measurements is likely the variation in thickness between different values of $\delta$.
In this article we have aimed to understand oxide barriers in JJs due to their technological importance in superconducting qubits. A summary of different thicknesses inferred within this paper is shown in Table 1. We fit the Simmons model to non-linear IVs to extract a single-valued thickness value to each junction $\sim$0.78 nm. We build Monte-Carlo protocols to simulate the IV properties of barriers with multi-valued thicknesses drawn from example statistical distributions. Here we find that the distribution of thicknesses present in the junction, results in IVs that, when fitted, would imply a single-value thickness smaller than the average thickness. This is because conductivity is exponentially dependent on the barrier thickness, so any thinner-than-average regions contribute more than similarly thicker-than-average regions. These simulations indicate a barrier that is $\sim$1 nm thick with a standard deviation of $\sim$0.2 nm, however using different barrier heights, we infer different thickness distributions. The rectangular potential barrier used in the Simmons model likely results in an under-estimate of thickness relative to more realistic, smoothly increasing barriers. By considering the Monte Carlo simulations relative to STEM measurements we constrain the barrier height to the range 0.8 eV $\lesssim\phi\lesssim$ 1.5 eV with best matching at $\sim$ 1.2 eV.